A certified public aacountant claims that more than 30% of all accountants advertise. A sample of 400 accounts in Rizal province showed that 112 had used some form of advertising. At a a=0.01,is there enough evidence to support the claim? (Use p-value method).

Solution:

Step 1:State the hypothesis.

H0:

H1:

Step 2:Compute for the value of one sample t test.zcomputed=

Step 3: Compute for th3 p-value.

Step 4:Decision rule:

Step 5: Conclusion


1
Expert's answer
2022-03-07T11:58:03-0500

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\le 0.3"

"H_1:p>0.3"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z: z > 2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{112/400-0.3}{\sqrt{\dfrac{0.3(1-0.3)}{400}}}=-0.87287"

Using the P-value approach:

The p-value is "p=P(Z>-0.87287)=0.808633," and since "p=0.808633>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is greater than "0.3," at the "\alpha = 0.01"significance level.